# PolyKAN: A Polyhedral Analysis Framework for Provable and Minimal KAN Compression

### v1

@@ 5.1 Algorithm Overview
[…]
10:return $\mathcal{N}^{\prime}$
The error allocation in Step 1 distributes the global error budget $\epsilon$ across layers according to the error propagation analysis, ensuring that the cumulative effect satisfies the global bound.
5.2 Optimal Spline Compression
The core of PolyKAN is the optimal compression of individual spline functions via dynamic programming.
Algorithm 2 OptimalSplineCompression
1:function OptimalSplineCompression($s$, $\delta$)
2:  $T\leftarrow[t_{0},t_{1},\ldots,t_{k}]$ $\triangleright$ Knot sequence of $s$
3:  $n\leftarrow\text{length}(T)$
4:  Initialize $dp[0\ldots n-1]$, $prev[0\ldots n-1]$ $\triangleright$ Dynamic programming arrays
5:  $dp[0]\leftarrow 1$, $prev[0]\leftarrow-1$ $\triangleright$ First knot always retained
6:  for $i\leftarrow 1$ to n1n-1 do
7:   $dp[i]\leftarrow\infty$
8:   for $j\leftarrow 0$ to i1i-1 do
9:     if $\text{CheckMergability}(T,j,i,\delta)$ then $\triangleright$ Test if regions between $j$ and $i$ can be merged
10:      $cost\leftarrow dp[j]+1$ $\triangleright$ Retain knot $i$
11:      if $cost<dp[i]$ then
12:        $dp[i]\leftarrow cost$, $prev[i]\leftarrow j$
13:      end if
14:     end if
15:   end for
16:  end for
17:  $new\_knots\leftarrow\text{BacktrackKnots}(prev,T)$
18:  $new\_coeffs\leftarrow\text{ComputeOptimalCoefficients}(T,new\_knots,\delta)$
19:  return $\text{ConstructSpline}(new\_knots,new\_coeffs)$
20:end function
The function $\text{CheckMergability}(T,j,i,\delta)$ verifies whether all regions between knots $j$ and $i$ can be merged into a single polynomial segment with approximation error at most $\delta$. This check can be implemented efficiently using polynomial approximation theory and interval analysis.
5.3 Theoretical Analysis of PolyKAN
We now establish the formal guarantees of the PolyKAN algorithm.
Theorem 3 (Correctness).
For any KAN network $\mathcal{N}$ and error tolerance $\epsilon>0$, the PolyKAN algorithm returns a compressed network $\mathcal{N}^{\prime}$ satisfying $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$.
@@ 5.3 Theoretical Analysis of PolyKAN
By construction, each spline compression in Algorithm 2 ensures that the local error is within the allocated budget $\delta_{l}$. The error allocation in Algorithm 1 ensures that the cumulative effect of these local errors satisfies the global bound $\epsilon$, following from the error propagation analysis in Lemma 4. ∎
Theorem 4 (Optimality).
For each individual spline function, Algorithm 2 finds the minimal number of knots that achieves the specified error bound $\delta$.
@@
The dynamic programming formulation in Algorithm 2 systematically explores all possible knot sequences. The cost $dp[i]$ represents the minimum number of knots needed to compress the spline up to knot $i$ while satisfying the error bound. The recurrence relation:
$dp[i]=\min_{j<i}\left\{dp[j]+1:\text{regions between $j$ and $i$ are $\delta$-mergable}\right\}$
ensures that all possible merging patterns are considered, guaranteeing optimality for each spline. ∎
Theorem 5 (Complexity).
For a KAN network with $L$ layers, at most $n$ nodes per layer, and at most $k$ knots per spline, the PolyKAN algorithm has time complexity $O(Ln^{2}k^{3})$.
Proof.
Algorithm 2 processes each spline in $O(k^{3})$ time: the outer loop runs $k$ times, the inner loop runs $k$ times, and each mergability check can be performed in $O(k)$ time using appropriate polynomial approximation techniques. With $O(Ln^{2})$ splines in total, the overall complexity is $O(Ln^{2}k^{3})$. ∎
These theoretical guarantees distinguish PolyKAN from heuristic compression methods, providing both performance bounds and formal correctness guarantees.
6 Theoretical Analysis and Case Study
6.1 Computational Complexity Analysis
The $O(Ln^{2}k^{3})$ complexity of PolyKAN is polynomial in all network parameters, making it computationally feasible for practical KAN architectures. The cubic dependence on $k$ is manageable since the number of knots per spline is typically small (10-100 in practice). Furthermore, the algorithm exhibits substantial parallelism, as different splines within and across layers can be compressed independently.
6.2 Error Bound Tightness
A crucial aspect of our theoretical analysis is establishing the tightness of our error bounds.
Proposition 1 (Error Bound Tightness).
There exist KAN networks and input distributions for which the error propagation bounds in Lemma 4 are tight. That is, for any $\delta>0$, there exists a KAN layer such that:
$\|\Phi-\tilde{\Phi}\|_{\infty}=n\cdot\delta$
when $\|s_{j,i}-\tilde{s}_{j,i}\|_{\infty}=\delta$ for all $i,j$.
Proof.
Consider a single KAN layer with $s_{j,i}(x)=c_{j,i}$ (constant functions) and inputs that simultaneously maximize all approximation errors. In this case, the errors accumulate linearly across dimensions, achieving the worst-case bound. ∎
This proposition demonstrates that our error analysis is not overly conservative and provides meaningful guarantees for practical compression scenarios.
6.3 Case Study: Illustrative Example
To elucidate the PolyKAN framework, we present a minimal case study compressing a univariate KAN approximating $f(x)=\sin(2\pi x)$ on $[0,1]$.
Original KAN: Single spline with knots at $[0,0.2,0.4,0.6,0.8,1.0]$ (5 regions).
Compression: With $\epsilon=0.1$, PolyKAN executes as follows:
  1. 1.
    Check mergability of adjacent regions: $0,0.2]\cup[0.2,0.4]$ is $\epsilon$-mergable.
  2. 2.
    Update knot sequence to $[0,0.4,0.6,0.8,1.0]$ (4 regions).
  3. 3.
    Check further merging: $0.4,0.6]\cup[0.6,0.8]$ is $\epsilon$-mergable.
  4. 4.
    Final knot sequence: $[0,0.4,0.8,1.0]$ (3 regions).
This example demonstrates how PolyKAN systematically reduces complexity while maintaining the error guarantee. The dynamic programming approach ensures that this is indeed the optimal compression for the given $\epsilon$.
7 Discussion and Conclusion
7.1 Theoretical Contributions
The PolyKAN framework represents a significant advance in the theory of neural network compression by providing:
  • Formal Guarantees: Unlike heuristic methods, PolyKAN provides provable bounds on both compression ratio and approximation error.
  • Optimality Certificates: The dynamic programming algorithm guarantees minimal compression for each spline function under the given error tolerance.
  • Computational Efficiency: Polynomial-time complexity makes the approach practical for realistic network sizes.
Our work bridges spline theory, polyhedral geometry, and algorithm design, establishing a new paradigm for trustworthy compression of interpretable neural architectures.
7.2 Limitations and Future Directions
While PolyKAN provides strong theoretical foundations, several limitations suggest fruitful directions for future research:
  • Layer-Wise Optimality: The current approach ensures optimal compression for each spline independently, but global optimality across all splines simultaneously remains an open challenge.
  • Error Allocation: Our error propagation bounds, while tight in worst-case, may be conservative for typical inputs. Input-aware error allocation could improve compression ratios.
  • Extended Spline Families: The theory currently focuses on B-splines; extension to other spline families and activation functions would broaden applicability.
Future work will explore these directions, with the goal of developing even more powerful compression frameworks with strong theoretical guarantees.
7.3 Conclusion
We have presented PolyKAN, a comprehensive theoretical framework for KAN compression with provable guarantees. By leveraging the intrinsic polyhedral structure of KANs, we have developed a complete theory of $\epsilon$-equivalent compression, designed an optimal algorithm with polynomial complexity, and established formal guarantees for correctness and minimality. This work provides a foundation for trustworthy compression of interpretable neural architectures, with potential applications in scientific computing, resource-constrained deployment, and theoretical computer science.

### v2

@@ 5.2 Approximate Compression for Multilayer KANs
9:return compressed network $\mathcal{N}^{\prime}$
Theorem 6 (Approximation Guarantee).
The compressed network $\mathcal{N}^{\prime}$ produced by Algorithm 2 satisfies $\|\mathcal{N}-\mathcal{N}^{\prime}\|_{\infty}\leq\epsilon$.
@@
By the single-layer error propagation lemma, each layer introduces at most $n_{l}\cdot(\epsilon_{l}/n_{l})=\epsilon_{l}$ error. The accumulation of errors across layers is controlled through error budget allocation, ensuring total error does not exceed $\sum_{l=1}^{L}\epsilon_{l}=\epsilon$. ∎
5.3 Optimality Gap Analysis
Although layered compression cannot guarantee global optimality, we can quantify its optimality gap.
Theorem 7 (Approximation Ratio).
Under the assumption of uniform error budget allocation, the gap between the compression ratio achieved by Algorithm 2 and that of the optimal solution is bounded polynomially by the network depth and width.
@@ 5.3 Optimality Gap Analysis
Let $OPT$ be the number of knots in the globally optimal compression and $ALG$ be the number obtained by the algorithm. Since each layer is compressed independently optimally and error propagation is linear, the optimality gap is bounded by the optimality of error budget allocation across layers. Specifically, there exists a constant $C$ (dependent on network structure) such that $ALG\leq C\cdot OPT$. ∎
6 Conclusion and Future Work
We have presented PolyKAN, a theoretical framework for KAN compression with provable guarantees. Although optimal KAN compression is NP-hard, we leveraged the axis-aligned structure of KANs to design efficient dynamic programming algorithms that guarantee optimality for univariate splines and provide approximation guarantees for the entire network.
Future work includes developing improved algorithms with better approximation ratios, investigating information-theoretic lower bounds for KAN compression, and extending the framework to other types of spline functions and network architectures.